An effective rate of 29,61% corresponds to a nominal rate, compounded weekly, of


[1] 34,35%.

[2] 29,53%.

[3] 26%.

[4] 29,61%.

[5] none of the above


1
Expert's answer
2021-05-24T18:05:09-0400

Let effective interest rate="i" ,


nominal interest rate is "j" ,


"i=100((\frac{1+j}{m})^m-1)"


"29.61=100((1+\frac {r} {52} )^{52}-1 )"


Put r=26,


"29.61=100((1+\frac {26} {52} )^{52}-1 )" ,


29.61=29.61.


LHS=RHS.